how can we find the value of sin22degree or all such angles?

Grade:12

4 Answers

Satyajit Samal
9 years ago
If you mean sin 22.5 degree, then we can find it using sin half angle formula.
\sin 22.5^{o}=\sin \frac{45}{2}^{o}= \sqrt {\frac{1- \cos 45^{o}}{2}}\\ = \sqrt {\frac{\sqrt{2}-1}{2 \sqrt{2}}}
sin 22 degree can not be found using this method. One way to find approximate value of sin 22 degree is to convert degree into radian and then use the sin x expansion formula:\sin x = x - \frac{x^3}{3!}+\frac{x^5}{5!}-\frac{x^7}{7!}+............
Hope that clears your doubt.
vaishali chandel
9 years ago
please give me more explanation for expansion of sin22 degree
 
vaishali chandel
9 years ago
please give me more explanation for expansion of sin22 degree
 
Kartik Sawant
8 years ago
Thanks to the expert, i got answer to another question.